# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ Appendix C Impact of Attackers’ Knowledge on Robustness: Comparison of Attack Settings
This appendix delves into the influence of varying levels of attackers’ knowledge about the stochastic components in diffusion processes on the robustness of diffusion-based models. We specifically assess the individual contributions of the forward and reverse diffusion processes to model robustness across different attack scenarios.
@@ C.1 Stochastic Elements in the Diffusion Processes
To elucidate the impact of the attacker’s knowledge, it is crucial to understand the stochastic elements integral to the diffusion processes, which are pivotal for the model’s robustness.
In the forward diffusion process, Gaussian noise is incorporated into the input data to derive a noisy version $x_{t}$:
$x_{t}=\sqrt{\bar{\alpha}_{t}}\,x+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{f},$ (10)
where $\epsilon_{f}\sim\mathcal{N}(0,I)$ is sampled once per input.
@@
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}t}},\bm{\epsilon}_{\bm{\theta}(x_{t},t)}\right)+\sigma_{t}\epsilon_{t},$ (11)
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is sampled at each reverse step. In contrast, for the Denoising Diffusion Implicit Model (DDIM), the reverse process is deterministic, and no noise $\{\epsilon_{t}\}_{t=1}^{T}$ is added.
@@ C.2 Attack Settings and Attacker Knowledge
We delineate four distinct attack scenarios, each characterized by the extent of information available to the attacker, particularly concerning the Gaussian noise variables in the diffusion processes. Table 10 provides a summary of the attacker’s knowledge in each scenario.
Table 10: Information accessible to the attacker in different attack settings. $\epsilon_{f}$ denotes the Gaussian noise in the forward process, and $\{\epsilon_{t}\}_{t=1}^{T}$ represents the Gaussian noise in the reverse process.
@@
Forward Process Noise $\epsilon_{f}$$\times$ $\times$
Reverse Process Noise $\{\epsilon_{t}\}_{t=1}^{T}$$\times$ $\times$
In the conventional white-box attack setting, the attacker possesses comprehensive knowledge of the model architecture and parameters but lacks insight into the stochastic elements used during inference ($\epsilon_{f}$ and $\{\epsilon_{t}\}_{t=1}^{T}$). The DWFwd setting grants the attacker knowledge of the Gaussian noise in the forward diffusion process ($\epsilon_{f}$). Conversely, the DWRev setting provides the attacker with knowledge of the Gaussian noise introduced during the reverse diffusion steps ($\{\epsilon_{t}\}_{t=1}^{T}$). The DWBoth setting offers the attacker complete access to all stochastic elements, $\epsilon_{f}$ and $\{\epsilon_{t}\}_{t=1}^{T}$. By manipulating the attacker’s knowledge in this manner, we isolate the individual effects of the forward and reverse diffusion processes on model robustness.
@@ C.3 Implications of the Attacker’s Knowledge of Stochastic Elements
The attacker’s capability to craft potent adversarial examples is significantly influenced by their knowledge of the stochastic elements in the diffusion processes. When these elements are unknown to the attacker, they must independently sample noise variables, leading to discrepancies between their approximations and the actual behavior of the victim model. Conversely, if the attacker is privy to the exact noise variables used during inference, they can precisely mimic the model’s behavior, markedly boosting the efficacy of their attack.
Attacker Without Knowledge of Stochastic Elements
In scenarios where the attacker lacks access to specific noise variables $\epsilon_{f}$ and $\{\epsilon_{t}\}_{t=1}^{T}$, the model’s output becomes unpredictable from the attacker’s viewpoint. The attacker must then optimize the expected value of the loss function over the distribution of these stochastic elements. The optimization problem for devising an adversarial example $x^{\text{adv}}$ is formulated as:
$x^{\text{adv}}=\arg\max_{\|x^{\text{adv}}-x\|\leq\delta}\,\mathbb{E}_{\epsilon_{f},\{\epsilon_{t}\}}\left[\mathcal{L}\left(f(x^{\text{adv}};\epsilon_{f},\{\epsilon_{t}\}),y\right)\right],$ (12)
where $\delta$ specifies the permissible perturbation magnitude, $\mathcal{L}$ is the loss function, $f$ represents the model’s output given the input and stochastic elements, and $y$ is the actual class label.
Attacker With Knowledge of Stochastic Elements
Should the attacker possess exact knowledge of the noise variables $\epsilon_{f}$ and $\{\epsilon_{t}\}_{t=1}^{T}$ utilized during the model’s inference, they can accurately emulate the victim classifier’s behavior. The stochastic processes become deterministic from the attacker’s perspective, facilitating the formulation of the optimization problem as:
$x^{\text{adv}}=\arg\max_{\|x^{\text{adv}}-x\|\leq\delta}\,\mathcal{L}\left(f(x^{\text{adv}};\epsilon_{f},\{\epsilon_{t}\}),y\right).$ (13)
@@ C.4 Effect of Attacker’s Knowledge on Model Robustness
Attack Setting Robust Accuracy ($l_{\infty}$)
@@
Conventional White-Box Attack
In this setting, the attacker fully understands the model’s architecture and parameters but lacks knowledge of the stochastic elements ($\epsilon_{f}$ and $\{\epsilon_{t}\}_{t=1}^{T}$) used during inference. The model’s output remains unpredictable due to the stochasticity of both diffusion processes, making it challenging for the attacker to generate effective adversarial examples (reaching robust accuracy of 47.27%).
DWFwd
Here, the attacker is aware of the Gaussian noise $\epsilon_{f}$ used in the forward diffusion process but not of the noise $\{\epsilon_{t}\}_{t=1}^{T}$ in the reverse process. This partial knowledge allows the attacker to accurately simulate the forward process, reducing uncertainty in this phase. However, the reverse process remains unpredictable. The slight decrease in robust accuracy to 45.41% suggests that while forward process stochasticity contributes to robustness, its effect is somewhat diminished when compromised.
DWRev
In this scenario, the attacker knows the noise variables $\{\epsilon_{t}\}_{t=1}^{T}$ used in the reverse diffusion steps but not the forward process noise $\epsilon_{f}$. This knowledge enables the attacker to align their strategy more closely with the actual behavior of the model during reverse diffusion, resulting in a more noticeable drop in robust accuracy to 35.25%. The reverse process’s stochasticity appears to play a more critical role in model robustness compared to the forward process.
DWBoth
When the attacker has comprehensive knowledge of both the forward and reverse process noise variables, they can replicate both diffusion processes accurately, eliminating any stochasticity from their perspective. This complete predictability allows for precise adversarial example crafting, leading to a significant reduction in robust accuracy to 16.80%. This demonstrates that the combined stochastic elements are crucial for maintaining robustness; when fully exposed, the model’s defense mechanisms are substantially weakened.
@@ Appendix D The Role of Stochasticity in DBP Compared to Certified Defense Methods
In this appendix section, we delve deeper into the role of randomness in Diffusion-Based Prediction (DBP) models and contrast it with its role in certified defense methods such as randomized smoothing (Cohen et al., 2019). While both approaches incorporate stochasticity, their mechanisms and implications for adversarial robustness differ significantly.
@@
  • Conventionally, the classification models discussed in the studies of adversarial robustness can be viewed as mappings from input space $X$ to the label space $Y$. However, DBP additionally involves a random variable $\epsilon\in E$ that determines the random sampling in the forward and reverse processes (which can be the random seed in implementation). Hence, a DBP model $f$ can be viewed as the mapping $f:(X,E)\rightarrow Y$.
@@
  • Previous studies on randomized smoothing treat the randomized model $f$ as a mapping $f:X\rightarrow P_{Y}$, where $P_{Y}$ is the space of label distribution. Typically, the final prediction can be formulated as $F(x)=\arg\max_{c}[f(x)]_{c}$, i.e., the class $c$ with the highest probability in the output distribution $\mathbb{f}(x)$. Apparently, $F$ deterministically maps $X$ to $Y$, consistent with the conventional models.
@@
  • Recent studies on DBP also regard the model as $f:X\rightarrow P_{Y}$, without explicitly studying the role of $\epsilon$. The key difference between DBP and randomized smoothing is that the final prediction for an input $x$ is directly sampled from the distribution $f(x)$ for once, instead of sampling multiple times to approximate $F(x)$ as in randomized smoothing.
@@
  • In this paper, we revisit DBP by treating the randomized model $f$ as the mapping $f:(X,E)\rightarrow Y$ and studying the role of $\epsilon\in E$ as an input of $f$. From this perspective, the conventional adversarial setting assuming full knowledge of the model parameters (but not $\epsilon$) is not a complete white box, which motivates us to study the DW-box setting.
@@
  • From our perspective, we can clearly point out the difference between DBP and randomized smoothing in terms of the loss landscape. Given an input $x_{0}$, the local loss landscape for a DBP model $f$ is not deterministic as it also depends on $\epsilon$. Although the expected loss landscape over $\epsilon\in E$ may be smooth, it does not suggest the robustness of DBP, as $\epsilon$ is fixed during a single inference run of DBP. Indeed, our study suggests that given $x_{0}$ and a fixed $\epsilon_{0}$, the local landscape of DBP is likely not smooth. In contrast, the loss landscape of a randomized smoothing model $F$ may be smooth as it is the average landscape over multiple $\epsilon$. To conclude, we argue that the random noise itself may not smooth the loss landscape, but the average over random noises may.
Appendix E Attack Method and Settings
@@
Previous assessments of DBP robustness have often utilized potentially unreliable methods. In particular, due to the iterative denoising process in diffusion models, some studies resort to mathematical approximations of gradients to reduce memory constraints (Athalye et al., 2018) or to circumvent the diffusion process during backpropagation (Wang et al., 2022). Furthermore, the reliability of AutoAttack, a widely used evaluation method, in assessing the robustness of DBP models is questionable. Although AutoAttack includes a Rand version designed for stochastic models, Nie et al. (2022) have found instances where the Rand version is less effective than the Standard version in evaluating DBP robustness.
@@
To improve the robustness evaluation of diffusion-based purification (DBP) models, we implement several modifications. First, to ensure the accuracy of the gradient computations, we compute the exact gradient of the entire diffusion classification pipeline. To mitigate the high memory requirements in diffusion iterative denoising steps, we use gradient checkpointing (Chen et al., 2016) techniques to optimize memory usage. In addition, to deal with the stochastic nature of the DBP process, we incorporate the Expectation over Transformation (EoT) method to average gradients across different attacks. We adopt EoT with 10 iterations, and a detailed discussion of the choice of EoT iterations can be found in Appendix A. We also use the Projected Gradient Descent (PGD) attack instead of AutoAttack for our evaluations [^1] We discover a bug in the Rand version of AutoAttack that causes it to overestimate the robustness of DBP. After fixing this, AutoAttack gives similar results to PGD attacks, but at a much higher computational cost. We discuss this in detail in Appendix L.. Our revised robustness evaluation revealed that DBP models, such as DiffPure and GDMP, perform worse than originally claimed. DiffPure’s accuracy dropped from a claimed 70.64% to an actual 55.96%, and GDMP’s from 90.10% to 40.97%. These results emphasize the urgent need for more accurate and reliable evaluation methods to properly assess the robustness of DBP models. Similar evaluation protocols are also applied in Chen et al. (2023); Kang et al. (2024).

### v3

@@ Appendix C Impact of Attackers’ Knowledge on Robustness: Comparison of Attack Settings
This section explores how the robustness of diffusion-based models is affected by the levels of knowledge that attackers possess concerning stochastic components in diffusion processes. We analyze how the knowledge of forward and reverse diffusion processes impacts model robustness in various attack settings.
@@ C.1 Stochastic Elements in the Diffusion Processes
Unlike deterministic models, where outputs are solely determined by the inputs, diffusion models introduce stochastic elements that also influence the outcomes. To explain the attacker’s knowledge, we first explain the stochastic components involved in diffusion processes.
In the forward diffusion process, Gaussian noise is incorporated into the input data to derive a noisy version $\bm{x}_{t}$:
$\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\,\bm{x}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon}_{f},$ (10)
where $\bm{\epsilon}_{f}\sim\mathcal{N}(\bm{0},\bm{I})$ is sampled once per input.
@@
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}t}}\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\right)+\sigma_{t}\bm{\epsilon}_{t},$ (11)
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,I)$ is sampled at each reverse step. In contrast, for the Denoising Diffusion Implicit Model (DDIM), the reverse process is deterministic, and no noise $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$ is added.
@@ C.2 Attack Settings and Attacker Knowledge
We define four attack settings based on the information accessible to the attacker, particularly regarding the Gaussian noise variables involved in the diffusion process. Table 10 provides a summary of the attacker’s knowledge in each setting.
Table 10: Information accessible to the attacker in different attack settings. $\bm{\epsilon}_{f}$ denotes the Gaussian noise in the forward process, and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$ represents the Gaussian noise in the reverse process.
@@
Forward Process Noise $\bm{\epsilon}_{f}$$\times$ $\times$
Reverse Process Noise $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$$\times$ $\times$
In the conventional white-box attack setting, the attacker possesses comprehensive knowledge of the model architecture and parameters but lacks insight into the stochastic elements used during inference ($\bm{\epsilon}_{f}$ and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$). The DWFwd setting grants the attacker knowledge of the Gaussian noise in the forward diffusion process ($\bm{\epsilon}_{f}$). Conversely, the DWRev setting provides the attacker with knowledge of the Gaussian noise introduced during the reverse diffusion steps ($\{\bm{\epsilon}_{t}\}_{t=1}^{T}$). The DWBoth setting offers the attacker complete access to all stochastic elements, $\bm{\epsilon}_{f}$ and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$. By manipulating the attacker’s knowledge in this manner, we isolate the individual effects of the forward and reverse diffusion processes on model robustness.
@@ C.3 Implications of the Attacker’s Knowledge of Stochastic Elements
The attacker’s capability to craft potent adversarial examples is significantly influenced by their knowledge of the stochastic elements in diffusion processes. When these elements are unknown to the attacker, they must independently sample noise variables, leading to discrepancies between their approximations and the actual behavior of the victim model. Conversely, if the attacker has access to the exact noise variables used during inference, they can accurately replicate the model’s behavior, significantly enhancing the effectiveness of their attack.
Attacker Without Knowledge of Stochastic Elements.
In scenarios where the attacker lacks access to the specific noise variables $\bm{\epsilon}_{f}$ and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$, the model’s output becomes unpredictable from the attacker’s perspective. Consequently, the attacker must optimize the expected value of the loss function over the distribution of these stochastic elements. The optimization problem for generating an adversarial example $\bm{x}^{\text{adv}}$ is formulated as:
$\bm{x}^{\text{adv}}=\operatorname{arg\,max}_{\|\bm{x}^{\text{adv}}-\bm{x}\|\leq\delta}\,\mathbb{E}_{\bm{\epsilon}_{f},\{\bm{\epsilon}_{t}\}_{t=1}^{T}}\left[L\left(f(\bm{x}^{\text{adv}};\bm{\epsilon}_{f},\{\bm{\epsilon}_{t}\}_{t=1}^{T}),y\right)\right],$ (12)
where $\delta$ specifies the permissible perturbation magnitude, $L$ is the loss function, $f$ represents the model’s output given the input and stochastic elements, and $y$ is the actual class label.
Attacker With Knowledge of Stochastic Elements.
If the attacker possesses precise knowledge of the noise variables $\bm{\epsilon}_{f}$ and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$ used during the model’s inference, they can accurately replicate the behavior of the victim classifier. In this case, the stochastic processes become deterministic from the attacker’s perspective, enabling the optimization problem to be formulated as:
$\bm{x}^{\text{adv}}=\operatorname{arg\,max}_{\|\bm{x}^{\text{adv}}-\bm{x}\|\leq\delta}\,L\left(f(\bm{x}^{\text{adv}};\bm{\epsilon}_{f},\{\bm{\epsilon}_{t}\}_{t=1}^{T}),y\right).$ (13)
@@ C.4 Effect of Attacker’s Knowledge on Model Robustness
Attack Setting Robust Accuracy ($\ell_{\infty}$)
@@
Conventional White-Box Attack.
In this setting, the attacker has full knowledge of the model’s architecture and parameters but lacks access to the stochastic components ($\bm{\epsilon}_{f}$ and $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$) involved during inference. Hence, the model’s output is unpredictable due to the stochasticity of both diffusion processes, making it challenging for the attacker to generate effective adversarial examples (reaching robust accuracy of 47.27%).
DWFwd.
Here, the attacker is aware of the Gaussian noise $\bm{\epsilon}_{f}$ used in the forward diffusion process but does not have knowledge of the noise $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$ used in the reverse process. This partial information enables the attacker to replicate the forward process, reducing uncertainty. However, the reverse process remains unpredictable. The exposure of the forward process leads to a slight decrease in robust accuracy to 45.41%.
DWRev.
In this case, the attacker knows the noise variables $\{\bm{\epsilon}_{t}\}_{t=1}^{T}$ used in the reverse diffusion process but lacks knowledge of the forward process noise $\bm{\epsilon}_{f}$. This enables the attacker to align their strategy more closely with the behavior of the model during the reverse diffusion phase, resulting in a more substantial drop in robust accuracy to 35.25%. These findings suggest that the stochasticity of the reverse process plays a more critical role in maintaining model robustness than that of the forward process.
DWBoth.
In this case, the attacker possesses full knowledge of the noise variables for both the forward and reverse diffusion processes, allowing them to precisely replicate both processes and eliminate stochasticity from their perspective. This complete predictability enables the attacker to craft highly effective adversarial examples, leading to a significant reduction in robust accuracy to 16.80%. These results underscore the importance of the stochastic elements in preserving model robustness; when fully exposed, the model’s defense mechanisms are substantially weakened.
@@ Appendix D The Role of Stochasticity in DBP Compared to Certified Defense Methods
In this appendix section, we delve deeper into the role of stochasticity in Diffusion-Based Prediction (DBP) models and contrast it with its role in certified defense methods such as randomized smoothing (Cohen et al., 2019). While both approaches incorporate stochasticity, their mechanisms and implications for adversarial robustness differ significantly.
@@
  • Conventionally, the classification models discussed in the studies of adversarial robustness can be viewed as mappings from input space $\mathcal{X}$ to the label space $\mathcal{Y}$. However, DBP additionally involves a random variable $\bm{\epsilon}\in\mathcal{E}$ that determines the random sampling in the forward and reverse processes (which can be the random seed in implementation). Hence, a DBP model protected classifier $f$ can be viewed as the mapping $f:(\mathcal{X},\mathcal{E})\rightarrow\mathcal{Y}$.
@@
  • Previous studies on randomized smoothing treat the randomized model $f$ as a mapping $f:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$, where $\mathcal{P}_{\mathcal{Y}}$ is the space of label distribution. Typically, the final prediction can be formulated as $F(\bm{x})=\operatorname{arg\,max}_{c}[f(\bm{x})]_{c}$, i.e., the class $c$ with the highest probability in the output distribution $\mathbb{f}(\bm{x})$. Apparently, $F$ deterministically maps $\mathcal{X}$ to $\mathcal{Y}$, consistent with the conventional models.
@@
  • Recent studies on DBP also regard the model as $f:\mathcal{X}\rightarrow\mathcal{P}_{\mathcal{Y}}$, without explicitly studying the role of $\bm{\epsilon}$. The key difference between DBP and randomized smoothing is that the final prediction for an input $\bm{x}$ is directly sampled from the distribution $f(\bm{x})$ for once, instead of sampling multiple times to approximate $F(\bm{x})$ as in randomized smoothing.
@@
  • In this paper, we revisit DBP by treating the randomized model $f$ as the mapping $f:\mathcal{X},\mathcal{E})\rightarrow\mathcal{Y}$ and studying the role of $\bm{\epsilon}\in\mathcal{E}$ as an input of $f$. From this perspective, the conventional adversarial setting assuming full knowledge of the model parameters (but not $\bm{\epsilon}$) is not a complete white box, which motivates us to study the DW-box setting.
@@
  • From our perspective, we can clearly point out the difference between DBP and randomized smoothing in terms of the loss landscape. Given an input $\bm{x}_{0}$, the local loss landscape for a DBP model $f$ is not deterministic as it also depends on $\bm{\epsilon}$. Although the expected loss landscape over $\bm{\epsilon}\in\mathcal{E}$ may be smooth, it does not suggest the robustness of DBP, as $\bm{\epsilon}$ is fixed during a single inference run of DBP. Indeed, our study suggests that given $\bm{x}_{0}$ and a fixed $\bm{\epsilon}_{0}$, the local landscape of DBP is likely not smooth. In contrast, the loss landscape of a randomized smoothing model $F$ may be smooth as it is the average landscape over multiple $\bm{\epsilon}$. To conclude, we argue that the random noise itself may not smooth the loss landscape, but the average over random noises may.
Appendix E Evaluation Settings
@@
Previous assessments of DBP robustness have often utilized potentially unreliable methods. In particular, due to the iterative denoising process in diffusion models, some studies resort to mathematical approximations of gradients to reduce memory constraints (Nie et al., 2022) or to circumvent the diffusion process during backpropagation (Wang et al., 2022). Furthermore, the reliability of AutoAttack, a widely used evaluation method, in assessing the robustness of DBP models is questionable. Although AutoAttack includes a Rand version designed for stochastic models, Nie et al. (2022) have found instances where the Rand version is less effective than the Standard version in evaluating DBP robustness.
@@
To improve the robustness evaluation of diffusion-based purification (DBP) models, we implement several modifications. First, to ensure the accuracy of the gradient computations, we compute the exact gradient of the entire diffusion classification pipeline. To mitigate the high memory requirements in diffusion iterative denoising steps, we use gradient checkpointing (Chen et al., 2016) techniques to optimize memory usage. In addition, to deal with the stochastic nature of the DBP process, we incorporate the Expectation over Transformation (EoT) method to average gradients across different attacks. We adopt EoT with 10 iterations, and a detailed discussion of the choice of EoT iterations can be found in Appendix A. We also use the Projected Gradient Descent (PGD) attack instead of AutoAttack for our evaluations. We discover a bug in the Rand version of AutoAttack that causes it to overestimate the robustness of DBP. After fixing this, AutoAttackFixed gives similar results to PGD attacks, but at a much higher computational cost. Our revised robustness evaluation revealed that DBP models, such as DiffPure and GDMP, perform worse than originally claimed. DiffPure’s accuracy dropped from a claimed 70.64% to an actual 55.96%, and GDMP’s from 90.10% to 40.97%. These results emphasize the urgent need for more accurate and reliable evaluation methods to properly assess the robustness of DBP models. Similar evaluation protocols are also applied in Chen et al. (2023); Kang et al. (2024).
E.1 Evaluation with Fixed AutoAttack
AutoAttack (Croce & Hein, 2020), an ensemble of White-box and Black-box attacks, is a popular benchmark for evaluating model robustness. It is used in RobustBench (Croce et al., 2020) to evaluate over 120 models. However, Nie et al. (2022) finds that the Rand version of AutoAttack, designed to evaluate stochastic defenses, sometimes yields higher accuracy than the Standard version that is intended for deterministic methods. Our comparison of AutoAttack and PGD20-EoT10 in Table 12 also shows that the Rand version of AutoAttack gives higher accuracy than the PGD20-EoT10 attack.
We attribute this to the sample selection of AutoAttack. As an ensemble of attack methods, AutoAttack selects the final adversarial sample from either the original input or the attack results. However, the original implementation neglects stochasticity and considers an adversarial sample to be sufficiently adversarial if it gives a false result in one evaluation. To fix this, we propose a 20-iteration evaluation and select the adversarial example with the lowest accuracy. The flawed code is in the official GitHub main branch, git version $a39220048b3c9f2cca9a4d3a54604793c68eca7e$, and specifically in lines #125, #129, #133-136, #157, #221-225, #227-228, #231 of the file ’autoattack/autoattack.py’. We will open-source our updated code and encourage future stochastic defense methods to be evaluated against the fixed code. The code can now be found at: https://github.com/Buntender/auto-attack.
Following the fix, robust accuracy under AutoAttackFixed decreases by up to 10 points, producing results comparable to our PGD20-EoT10 test outcomes. However, using AutoAttack on DPDDPM took nearly 25 hours, five times longer than PGD20-EoT10. Therefore, we will use PGD20-EoT10 for most robustness evaluation.
DBP method $\ell_{\infty}$ $\ell_{2}$
DiffPure 62.11 56.25 55.96 81.84 76.37 75.78
DPDDPM 57.81 46.88 48.63 71.68 71.09 72.27
DPDDIM 50.20 40.62 44.73 77.15 70.70 71.68
Table 13: Clean and robust accuracy (%) on different DBP methods for CIFAR-10, evaluated with AutoAttackFixed (Rand version). All methods show consistent improvement when fine-tuned with ADDT.
DBP method Vanilla ADDT
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$
DiffPure 89.26 56.25 76.37 89.94 58.20 77.34
DPDDPM 85.94 46.88 71.09 85.64 48.63 72.27
DPDDIM 88.38 40.62 70.70 88.77 44.73 71.68